OpenStudy (anonymous):

algebra 0.01 help anyone? see pic

OpenStudy (anonymous):

is it some kind of cross multiplication or something?

OpenStudy (anonymous):

how do you get from the first bit to teh second bit is my question

OpenStudy (anonymous):

First multiply by 64

OpenStudy (anonymous):

\[540/512 a = 9a ^{4}\]

OpenStudy (anonymous):

Sometimes it is not necessary to simplify, do it later 9a^4=[9(64)a]/8 Next you divide both sides by a and by 9

OpenStudy (anonymous):

a cubed = 64/8 i guess that was easier than i thought

OpenStudy (anonymous):

It is just a little algebra, as in solving for x. You get x (or a) on one side and everything else on the other. A final step from your handwritten note is take the cube root of each side a^3=8 a =2

OpenStudy (anonymous):

oh i knew what to call it, just couldn't see how it was done in one step... thanks